Given income is P15.00 Px is 2.00/kg and Py is P1.00/pack
1
Expert's answer
2021-10-02T10:48:27-0400

There is no utility function equation, there is no way to calculate utility maximization taking into account the budget constraint.

The given data make it possible to determine what the same quantity of goods x and y can be purchased taking into account the reduced income.

"x=y=\frac {P}{Px+Py}=\frac{15}{2+1}=\frac{15}{3}=5"